According to the Beer-Lambert Law, the absorbance of a solution varies jointly as the concentration of solute and the length through which the light passes. A solution of albumin with a concentration of 11.64 μg/mL placed in a 1-cm wide cuvette has an absorbance of 0.121. What is the absorbance of a diluted solution of albumin with a concentration of 7.29 μg/mL?
Answer by Boreal(15235)

A=k*s*l
find k
0.121=k*11.64*1
k=0.121/11.64
=0.0104
Use in second part.
A=0.0104*7.29=
0.0758, assuming the same 1 cm wide cuvette.